# Estimates for Dirichlet Eigenvalues of the Schrödinger operator with the Kronig-Penney Model

### v1

@@ 3 Estimates for the Small Eigenvalues
where
$a_{r,k,n}(\lambda)=\sum_{n_{1},n_{2},...,n_{k}=-r}^{r}\frac{C_{n_{1}}C_{n_{2}}\cdots C_{n_{k}}(C_{n_{1}+n_{2}+\cdots+n_{k}}-C_{n_{1}+n_{2}+\cdots+n_{k}+2n})}{[\lambda-(n+n_{1})^{2}]\cdots[\lambda-(n+n_{1}+\cdots+n_{k})^{2}]}.$ (54)
Define the function
$K_{n}(\lambda):=\lambda-n^{2}-g_{n}(\lambda),$ (55)
where
$g_{n}(\lambda)=-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda).$ (56)
@@
Proof. First we prove (58) by using the mean-value theorem. To do this, we estimate $|g_{n}^{\prime}(\lambda)|=|\dfrac{d}{d\lambda}g_{n}(\lambda)|$. By (56), we have
@@
First let us estimate the summands of the term $|\dfrac{d}{d\lambda}a_{r,k,n}(\lambda)|$:
@@
For the estimation of the right-hand side of (62), we obtain
@@
$\qquad=\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)]}+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}$ (63)
Thus, by (62) and (63), we obtain
@@
where $g_{n}(x)$ is defined by (56). Since
$|g_{n}(\lambda_{n})|=\biggl|-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|$
$\qquad\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(\lambda_{n})|,$ (67)
@@
$|g_{n}(n^{2})|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(n^{2})|$
$\qquad<\frac{(b-a)}{2\pi n}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)}+\sum\limits_{k=2}^{r}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1)^{k}}$
@@
Proof. Without loss of generality we can take $x_{n,0}=n^{2}$. By (57), (58), and (66) we have
@@
Therefore it is enough to estimate $|x_{n,0}-\rho_{n}|$. By definitions of $\rho_{n}$ and $x_{n,0}$ we obtain
@@
and by the mean value theorem there exists $x\in[n^{2}-M,n^{2}+M]$ such that
@@
$\lambda_{1}$ $=0.984205232093$
$\lambda_{2}$ $=4.003110832419$
$\lambda_{3}$ $=9.013023482675$
$\lambda_{4}$ $=16.018415152245$
$\lambda_{5}$ $=25.010308870985$
$\lambda_{6}$ $=36.010794654577.$
@@
Declarations
  • Funding
    Not applicable.
  • Conflict of interest
    The author states that there is no conflict of interest.
  • Ethics approval
    Not applicable.
  • Consent to participate
    Not applicable.
  • Consent for publication
    Not applicable.
  • Availability of data and materials
    Not applicable.

### v2

@@ 3 Estimates for the Small Eigenvalues
where $a_{r,k,n}(\lambda)$ is obtained from (5) by taking finite sums instead of the infinite series, namely,
$a_{r,k,n}(\lambda):=\sum_{n_{1},n_{2},...,n_{k}=-r}^{r}\frac{C_{n_{1}}C_{n_{2}}\cdots C_{n_{k}}(C_{n_{1}+n_{2}+\cdots+n_{k}}-C_{n_{1}+n_{2}+\cdots+n_{k}+2n})}{[\lambda-(n+n_{1})^{2}]\cdots[\lambda-(n+n_{1}+\cdots+n_{k})^{2}]}.$
Define the functions
$K_{n}(\lambda):=\lambda-n^{2}-g_{n}(\lambda),\qquad g_{n}(\lambda)=-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda).$ (46)
@@
Proof. First, we prove (48) by using the mean-value theorem. To do this, we estimate $|g_{n}^{\prime}(\lambda)|=|\dfrac{d}{d\lambda}g_{n}(\lambda)|$. By (46), we have
@@
By estimating the summands of the term $|\dfrac{d}{d\lambda}a_{r,k,n}(\lambda)|$, we write
@@
For the estimation of the right-hand side of (52), we have
@@
$\qquad=\frac{(b-a)^{s+2}}{2M\sqrt{2}^{s}\pi^{s+1}(2n-1-M)^{s}[\sqrt{2}\pi(2n-1-M)-(b-a)]}+\frac{8(b-a)^{2}}{\pi^{2}(r+1)^{2}[(r+1)|r+1-2n|-M]}.$ (53)
Thus, by (52) and (53), we have
@@
where $g_{n}(x)$ is defined in (46). Since
$|g_{n}(\lambda_{n})|=\biggl|-C_{2n}+\sum\limits_{k=1}^{s}a_{r,k,n}(\lambda_{n})\biggr|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(\lambda_{n})|,$ (57)
@@
$|g_{n}(n^{2})|\leq|C_{2n}|+\sum\limits_{k=1}^{s}|a_{r,k,n}(n^{2})|<\frac{(b-a)}{2\pi n}+\frac{(b-a)^{2}}{M\pi^{2}(2n-1)}+\sum\limits_{k=2}^{r}\frac{(b-a)^{k+1}}{2M\sqrt{2}^{k}\pi^{k+1}(2n-1)^{k}}$
@@
Proof. Without loss of generality, we can take $x_{n,0}=n^{2}$. By (47), (48), and (56), we have
@@
Therefore, it is enough to estimate $|x_{n,0}-\rho_{n}|$. By definitions of $\rho_{n}$ and $x_{n,0}$, we obtain
@@
and by the mean value theorem, there exists $x\in[n^{2}-M,n^{2}+M]$ such that
@@
$\lambda_{1}$ $=0.984205232093,\quad\lambda_{2}=4.003110832419,\quad\lambda_{3}=9.013023482675$
$\lambda_{4}$ $=16.018415152245,\quad\lambda_{5}=25.010308870985,\quad\lambda_{6}=36.010794654577.$